In the laboratory you dissolve 23.7 g of iron(II) acetate in a volumetric flask and add water to a total volume of 375 mL.

What is the molarity of the solution?  M.

What is the concentration of the iron(II) cation?  M.

What is the concentration of the acetate anion?  M.
Answer #1

1)

Molar mass of Fe(CH3CO2)2,
MM = 1*MM(Fe) + 4*MM(C) + 6*MM(H) + 4*MM(O)
= 1*55.85 + 4*12.01 + 6*1.008 + 4*16.0
= 173.938 g/mol


mass(Fe(CH3CO2)2)= 23.7 g

use:
number of mol of Fe(CH3CO2)2,
n = mass of Fe(CH3CO2)2/molar mass of Fe(CH3CO2)2
=(23.7 g)/(1.739*10^2 g/mol)
= 0.1363 mol
volume , V = 3.75*10^2 mL
= 0.375 L


use:
Molarity,
M = number of mol / volume in L
= 0.1363/0.375
= 0.3633 M
Answer: 0.363 M


2)
[Fe2+] = [Fe(CH3CO2)2]
= = 0.3633 M
Answer: 0.363 M

3)
[CH3CO2-] = 2*[Fe(CH3CO2)2]
= = 2*0.3633 M
Answer: 0.727 M